The monthly salaries of contract workers in a small Jamaican shipping company are approximately normally distributed with a mean of $60,000 and a standard deviation of $4,000. What percent of people earn less than $55,000?


1
Expert's answer
2022-01-24T14:58:57-0500

Let "X="the monthly salary: "X\sim N(\mu, \sigma^2 )."

Given "\mu=\$60000, \sigma=\$4000."


"P(X<55000)=P(Z<\dfrac{55000-60000}{4000})"

"=P(Z<-1.25)=0.10565"

"10.565\%"